# Observational Auditing of Label Privacy

### v1

@@ 1 Introduction
Differential privacy (DP) auditing has become an important tool for evaluating privacy guarantees in machine learning systems. Recent advances in auditing methods that require only a single run have made it feasible to evaluate privacy for large-scale models without prohibitive computational costs (Steinke et al., 2024; Mahloujifar et al., 2025b). However, most existing auditing approaches still require modifying the training dataset by injecting known entropy or canary data, which limits their applicability in industry-scale environments, where modifications to the training data pipeline require significant engineering overhead.
@@
In this work, we propose a novel auditing methodology that eliminates the need for dataset modification. Our approach enables privacy evaluation using the natural nondeterminism present in the data distribution itself. We formalize and empirically validate this methodology in the setting of auditing Label DP, generalizable to privacy guarantees for any protected attribute. This capability addresses a significant gap in current auditing techniques, which have primarily focused on membership inference attacks (Shokri et al., 2017; Carlini et al., 2022) rather than attribute inference. In particular, existing methods for auditing Label DP either require adding out-of-distribution canaries to the training set (Malek et al., 2021) or are applicable only to a limited set of mechanisms (Busa-Fekete et al., 2024).
@@
Observational privacy auditing cannot be done unconditionally, without making certain assumptions about the underlying distribution (Hernán and Robins, 2020, Chapter 3). Unlike anecdotal instances of privacy violations (Barbaro and Zeller Jr, 2006; Narayanan and Shmatikov, 2008; Carlini et al., 2021), auditing seeks to provide statistically valid measurements of memorization. In other words, the objective of privacy auditing is to establish causality—demonstrating that a model behaves in a certain way because it was trained on specific data. Traditionally, most rigorous membership inference attacks establish and measure causal effects through randomized control trials (RCTs), which require interference with the training data. By reframing auditing as a security game between two adversarial parties, our approach eliminates the need for training-time intervention.
@@
Our key assumption is the availability of a distribution that approximates the ground truth. Concretely, the Label DP auditing mechanism relies on access to a proxy label-generating distribution. The proxy does not need to match the ground truth distribution, provided the adversary cannot distinguish between them (with reasonable computational resources). Under this assumption, the counterfactual examples generated by the challenger can be used to evaluate the attack on the model’s claimed Label DP guarantees without training data intervention. The attack can be made practical by using a model other than the target model as the proxy distribution. Further, in an incremental learning setting, earlier model checkpoints can be used as the proxy distribution, thus requiring no additional model training and minimal engineering overhead. The
@@
We demonstrate that our observational auditing framework provides accurate privacy bounds that match those obtained by interventional methods and provable privacy guarantees. Through a series of cryptography-inspired games we establish the theoretical foundations for auditing privacy without training data manipulation. By lowering the complexity of privacy auditing, our approach enables its application in a wider variety of contexts.
@@ 2 Background and prior work
Differential Privacy (DP) introduced by Dwork et al. (2006) is a leading framework for providing rigorous privacy guarantees in statistical data analysis and machine learning. In its standard formulation, DP bounds the impact that any single individual’s record has on the outcome of a computation by constraining how much the distribution of outputs may differ between neighboring datasets. The definition’s strong theoretical guarantees, resilience to arbitrary auxiliary data, and compositional properties have driven its adoption in academic research and industry deployments (Fioretto and Hentenryck, 2025).
The general DP framework can be adapted to settings where only certain parts of a dataset are considered private. This paper focuses on Label DP, which has emerged as an important objective for PPML, particularly in the domain of recommendation systems (Chaudhuri and Hsu, 2011; Ghazi et al., 2021; Malek et al., 2021; Wu et al., 2023). The following factors motivate Label DP as a uniquely valuable privacy concept:
@@
Complementing the strong worst-case guarantees of DP that bound the privacy loss from above on all inputs, privacy auditing empirically measures the privacy loss on concrete instances, providing a lower bound on DP’s numerical parameters. Privacy auditing can be used for finding bugs in claimed implementations of DP algorithms (Ding et al., 2018), advancing understanding of complex DP mechanisms (Malek et al., 2021; Nasr et al., 2023, 2025), or guardrailing models in a production environment (Agrawal and Book, 2025).
@@
Privacy auditing consists of two components: a privacy game between the challenger and the attacker, and an auditing analysis that translates the attacker’s success into lower bounds on the $({\epsilon},\delta)$-DP guarantee (or other forms of DP). The privacy game is characterized by the capabilities and resources of the parties, and the attacker’s goals, such as reconstruction, membership or attribute inference. Auditing analyses, typically used for membership inference, can be applied to any stochastic privacy game (Swanberg et al., 2025). We show theoretical results for our label inference attack building on Steinke et al. (2024) and Mahloujifar et al. (2025b).
@@
Membership inference attacks (MIAs)—where an adversary uses model access and knowledge of the data distribution to determine whether a sample was part of training—have received significant attention in the literature (Shokri et al., 2017; Yeom et al., 2018; Salem et al., 2019; Sablayrolles et al., 2019; Song and Shmatikov, 2019; Nasr et al., 2019; Leino and Fredrikson, 2020; Carlini et al., 2022; Ye et al., 2022; Zarifzadeh et al., 2024; Bertran et al., 2023). This attack category directly maps to the differential privacy guarantee where two neighboring datasets differ in the presence of one training sample.
@@
A less studied category is attribute inference attacks (Yeom et al., 2018) where the adversary reconstructs a protected attribute given access to a partial record, of which label inference attacks are a special case (Malek et al., 2021; Busa-Fekete et al., 2024). A difficulty and common pitfall with such attacks is to properly account for the adversary’s baseline success, achieved by exploiting knowledge of the data distribution and correlations between the public and protected attributes (Jayaraman and Evans, 2022). The label inference attack of Malek et al. (2021) uses canaries with random binary labels, which sets the adversary’s baseline accuracy to $0.5$ and allows Label DP to be audited via standard MIA analyses. However, this attack interferes with the training dataset and can affect model performance if too many out-of-distribution samples are injected.
@@
Most auditing methods in the literature are “interventional,” as their privacy game involves manipulating the training dataset: MIAs require excluding a subset of the data from training (Steinke et al., 2024; Mahloujifar et al., 2025b), whereas Malek et al. (2021) modifies the training labels. These requirements severely restrict the applications of these auditing methods. Instead, our label inference attack can run entirely post-training. MIAs can be stated as observational privacy games if the challenger is able to sample fresh samples from the distribution (Ye et al., 2022). However, obtaining new samples from the distribution (or from its close approximation), without affecting the training pipeline, remains an extremely challenging open problem (Meeus et al., 2025).
@@
Kazmi et al. (2024) propose an observational MIA that uses generative models such as generative adversarial networks (GANs) to obtain fresh samples. Differently from Kazmi et al. (2024), we propose a more general observational auditing framework (with observational MIA and LIA as special cases), and provide theoretical lower bounds on the privacy guarantees for these general attacks. A challenge with using synthetic samples is to account for how the shift between the generative model and the true distribution boosts the success of the attacker. Counterfactual labels are much easier to generate than entire samples (models are trained to do precisely this), and we can assume that an adversary with limited computation resources cannot distinguish between a label from the distribution and a label predicted from the best model the adversary can train.
@@
The closest observational label inference attack to ours is the recent work of Busa-Fekete et al. (2024), which measures the label reconstruction advantage of the adversary with and without access to the model. This metric is not translated into a lower bound on $\epsilon$} in the differential privacy guarantee; in fact, such translation would be difficult because the adversary has a different baseline success (prior) for each sample. Additionally, this approach requires estimating probabilities of the mechanism’s output given a particular label, limiting its applicability to simple mechanisms like randomized response and random label aggregation. In contrast, our label inference attack is observational and applicable to all mechanisms. It can audit label privacy in a statistically valid manner because it sets up a game where the baseline accuracy of the adversary is $0.5$ for all samples.
@@ 3 Preliminaries
In this work, we audit the simulation-based definition of differential privacy. It generalizes the traditional add/remove (or “leave-one-out”, or “zero-out”) notion of DP to support a privacy unit that is a subset of the sample’s attributes. The definition compares the distribution of a mechanism $M$ on dataset $D$ with that of a simulator that emulates the output of $M$ on $D$ without seeing the protected attributes of a record. See Appendix 8 for further discussion of this definition.
Definition 1 (Simulation-based privacy for protected attributes (Mahloujifar et al., 2025a)).
@@
For the more advanced notion of $f$-DP (Dong et al., 2020), a mechanism is f-SIM-DP if
$M(D)\approx_{f}\mathrm{Sim}(D^{\prime},x).$ (2)
We also say $M$ is $(\epsilon,\delta)$-SIM-DP (resp. f-SIM-DP) if there exists a simulator $\mathrm{Sim}$ for which (1) (resp. (2)) holds.
@@
An audit procedure takes the description of a mechanism $M$, a trade-off function $f$, a simulator $\mathrm{Sim}$ and decides whether the mechanism satisfies f-SIM-DP. We define it as a two-step process.
  • $\mathrm{Game}\colon\mathcal{M}\times\mathcal{S}\to\mathcal{O}$. The auditor runs a potentially randomized experiment/game using the description of mechanism $M\in\mathcal{M}$ and the simulator $\mathrm{Sim}$. The auditor receives the game output $o\in\mathcal{O}$.
  • $\mathrm{Evaluate}\colon\mathcal{O}\times\mathcal{F}\to\{0,1\}$. The output is 0 if the auditor rejects the hypothesis that $M$ satisfies f-SIM-DP based on evidence $o$, and $1$ otherwise.
@@ 4 Observational versus interventional privacy games
An observational privacy game considers the training dataset as a given. In contrast, an interventional privacy game interferes with the training data pipeline and the eventual dataset used for training. In this section, we formalize our observational privacy auditing framework. To that end, we introduce a generic attack game, generalizing Swanberg et al. (2025), as Algorithm 1.
The privacy game occurs between two parties: a challenger and an adversary. A key difference from Swanberg et al. (2025) is that we split the challenger algorithm into two stages: determining the training data $(G_{\text{intervention}})$ and determining the additional outputs provided to the adversary ($G_{\text{hint}}$). With this, we can separate observational games from interventional ones. Further we distinguish between the training dataset $D$ and additional game artifacts $S$, which are used by the challenger to set up a stochastic game (e.g., sampling random bits).
Algorithm 1 Generic Attack Game (adapted from Swanberg et al. (2025))
@@
The algorithm $G_{\text{intervention}}$ determines the training dataset for $M$, obtained from a potential modification of the fixed training set $D$. For instance, in the one-run MIA (Steinke et al., 2024) the artifacts are $S=(S_{i}\sim\mathsf{Bernoulli}(0.5)\colon i\in[m])$, where $m$ is the number of canaries. The training dataset $\overline{D}$ is obtained from $D$ by including all samples $x_{i}\in D$ where $S_{i}=1$ and excluding all samples where $S_{i}=0$.
The role of $G_{\text{hint}}$ is to collect additional information the challenger provides to the adversary, based on game artifacts, training data, and the output of the trained model $M(\overline{D})$. In the one-run MIA, $o_{2}$ is the vector of targets $x_{1},\dots,x_{m}$ from the training set $D$.
@@
Observational versus interventional MIA. We have described how the (interventional) one-run MIA (Steinke et al. (2024)) can be framed as Algorithm 1. We now present an observational one-run MIA (following Ye et al. (2022)).
@@
From the adversary’s perspective, the observational one-run MIA has the same distribution as the interventional MIA. For the challenger, the difference matters as the observational game does not alter the training pipeline. In this game, the source of counterfactual samples $x_{i}^{\prime}$ can be a distribution that approximates $\mathcal{D}$ sufficiently well, as is the case with the attack of Kazmi et al. (2024) who use generative models to sample $x_{i}^{\prime}$.

### v2

@@ 1 Introduction
Differential privacy (DP) auditing has become an important tool for evaluating privacy guarantees in machine learning systems. Recent advances in auditing methods that require only a single run have made it feasible to evaluate privacy for large-scale models without prohibitive computational costs (steinke2024privacy; mahloujifar2025auditing). However, most existing auditing approaches still require modifying the training dataset by injecting known entropy or canary data, which limits their applicability in industry-scale environments, where modifications to the training data pipeline require significant engineering overhead.
@@
In this work, we propose a novel auditing methodology that eliminates the need for dataset modification. Our approach enables privacy evaluation using the natural nondeterminism present in the data distribution itself. We formalize and empirically validate this methodology in the setting of auditing Label DP, generalizable to privacy guarantees for any protected attribute. This capability addresses a significant gap in current auditing techniques, which have primarily focused on membership inference attacks (shokri2017membership; carlini2022membership) rather than attribute inference. In particular, existing methods for auditing Label DP either require adding out-of-distribution canaries to the training set (meta-labeldp) or apply only to a limited set of mechanisms (fekete2024auditing).
@@
Observational privacy auditing cannot be done unconditionally, without making certain assumptions about the underlying distribution (hernan2020causal, Chapter 3). Unlike anecdotal instances of privacy violations (barbaro2006aol; narayanan2008-netflix; carlini2021extracting), auditing seeks to provide statistically valid measurements of memorization. In other words, the objective of privacy auditing is to establish causality—demonstrating that a model behaves in a certain way because it was trained on specific data. Traditionally, most rigorous membership inference attacks establish and measure causal effects through randomized control trials (RCTs), which require interference with the training data (MIA_cannot_prove25). By framing auditing as a security game between two adversarial parties, our approach eliminates the need for training-time intervention.
@@
Our key assumption is the availability of a distribution that approximates the ground truth data distribution. Concretely, the Label DP auditing mechanism relies on access to a proxy label-generating distribution. The proxy does not need to match the ground truth distribution, provided the adversary cannot distinguish between them (with reasonable computational resources). Under this assumption, the counterfactual examples generated by the challenger can be used to evaluate the attack on the model’s claimed Label DP guarantees without training data intervention. The attack can be made practical by using a model other than the target model as the proxy distribution. Further, in an incremental learning setting, earlier model checkpoints can be used as the proxy distribution, thus requiring no additional model training and minimal engineering overhead.
@@
We demonstrate that our observational auditing framework provides accurate privacy bounds that match those obtained by interventional methods. Through a series of cryptography-inspired games we establish the theoretical foundations for auditing privacy without training data manipulation. By lowering the complexity of privacy auditing, our approach enables its application in a wider variety of contexts.
Our approach and key contributions. In the rest of this section we focus on the high-level intuition and concepts, deferring the formal treatment for later. The standard definition of differential privacy is formulated as a bound on the stability in distribution of a randomized algorithm in response to small changes in its input, such as inclusion or exclusion of a single user’s data. Consequently, the auditing of differentially private implementations is typically presented as empirically testing these bounds by detecting changes in the output distribution on close inputs.
Definitions equivalent or closely related to differential privacy can be framed using the notion of a simulator whose role is to “create the illusion of stability” while protecting individual records in its input (CDP09; mahloujifar2025machine). In this approach, the privacy definition is operationalized as a game played between the simulator (“the defender”) and the adversary (“the attacker”), whose probability of success under the rules of this game is constrained by the privacy parameters.
Our first contribution is incorporating the simulator implementation into the auditing procedure. Simulation-based definitions are standard (indeed foundational) in theoretical cryptography (GM82; GMR85). The key distinction is that the cryptographic simulator is always a thought experiment, present in the proof but never implemented in practice (except for didactic purposes). In our approach, the auditor actually tests the simulator’s performance vis-à-vis the distinguisher in order to derive privacy loss estimates.
The main advantage of simulation-based DP definitions is that they are inherently compatible with observational auditing. Whereas the counterfactuals of the standard DP formulation are produced by the same process as the real-world execution, the simulator can be implemented quite differently. The other advantage of simulation-based DP, as used in mahloujifar2025machine, is the extension of the add/remove DP framework (rather than replacement) to the privacy of partial records.
Auditing simulation-based differential privacy (DP) differs significantly from the standard approach. Instead of evaluating privacy by having a (fixed, domain-agnostic) challenger evaluate the attacker’s success in breaching privacy guarantees, the process involves two competing algorithms: the simulator and the attacker. The simulator’s goal is to generate outputs that closely resemble those of the real mechanism, but without access to certain masked data, thereby confusing the attacker. In this framework, the attacker may succeed for two main reasons: (1) the audited mechanism genuinely leaks private data, or (2) the simulator fails to effectively mimic the real mechanism. Consequently, the audit may underestimate privacy loss if the attacker is not sufficiently strong, or overestimate it if the simulator’s outputs allow the attacker to easily distinguish between simulated and real data.
These trade-offs are not unique to our approach. In most cases, MIAs implicitly assume that the challenger can sample both members and non-members of the training set. Unless this capability is established before training the target model, typically through some form of intervention, the challenger’s options are limited, and none are airtight (MIA_cannot_prove25). By explicitly modeling the simulator as a participant in the privacy game, we allow the simulator’s output to diverge from the true data distribution and can quantify the resulting slack in auditing guarantees.
A setting where this approach is especially effective is in auditing Label DP, which evaluates the extent to which classifiers memorize sensitive labels. In this context, the simulator’s task is reduced to generating labels that closely match the true distribution. This task is not only fundamentally simpler than synthesizing complete data samples, but it also aligns directly with the objective of the target model. Thus, the architecture of the target model can be readily adapted for use by the simulator.
Our theoretical analysis builds upon the one-run auditing framework of steinke2024privacy and mahloujifar2025auditing. The novel part of the argument incorporates an additional parameter $\tau$ that controls the fidelity of the simulator into tail bounds on the adversary’s advantage as a function of privacy parameters. We demonstrate that the audit remains accurate as long as we have a high quality estimator of the true data distribution.
To summarize, in this work we:
  • Introduce a privacy auditing framework that does not interfere with model training. Instead, it runs a post-training game between a simulator that produces synthetic records (or record attributes) and an attacker that distinguishes between training and synthetic records.
  • Generalize differential privacy auditing guarantees from mahloujifar2025auditing to account for the gap between the simulator and the true data distribution.
  • Instantiate our framework for the case of auditing Label DP, giving an easy to implement attack in production settings, where labels are synthesized using earlier model checkpoints. We evaluate this attack empirically and show that it captures similar levels of memorization as interventional MIAs.
@@ 2 Background and prior work
Differential Privacy (DP) introduced by DMNS06 is a leading framework for providing rigorous privacy guarantees in statistical data analysis and machine learning. In its standard formulation, DP bounds the impact that any single individual’s record has on the outcome of a computation by constraining how much the distribution of outputs may differ between neighboring datasets. The definition’s strong theoretical guarantees, resilience to arbitrary auxiliary data, and compositional properties have driven its adoption in academic research and industry deployments (Fioretto2025).
The general DP framework can be adapted to settings where only certain parts of a dataset are considered private. This paper focuses on Label DP, which has emerged as an important objective for privacy preserving machine learning, particularly in the domain of recommendation systems (chaudhuri2011-labeldp; google-labeldp; meta-labeldp; wu2022does). The following factors motivate Label DP as a uniquely valuable privacy concept:
@@
Complementing the strong worst-case guarantees of DP that bound the privacy loss from above on all inputs, privacy auditing empirically measures the privacy loss on concrete instances, providing a lower bound on DP’s numerical parameters. Privacy auditing can be used for finding bugs in claimed implementations of DP algorithms (ding_etal-detecting_violations2018), advancing understanding of complex DP mechanisms (meta-labeldp; nasr_etal_tight_auditing23; nasr_etal_last_iterate25), or guardrailing models in a production environment (snap2025).
@@
Privacy auditing consists of two components: a privacy game between the challenger and the attacker, and an auditing analysis that translates the attacker’s success into lower bounds on the $({\epsilon},\delta)$-DP guarantee (or other forms of DP). The privacy game is characterized by the capabilities and resources of the parties, and the attacker’s goals, such as reconstruction, membership or attribute inference. Auditing analyses, typically used for membership inference, can be applied to any stochastic privacy game (swanberg2025). We show theoretical results for our label inference attack building on steinke2024privacy and mahloujifar2025auditing.
@@
Membership inference attacks (MIAs)—where an adversary uses model access and knowledge of the data distribution to determine whether a sample was part of training—have received significant attention in the literature (shokri2017membership; yeom2018privacy; salem2018mlleaks; sablayrolles2019white; song2019membership; nasr2019comprehensive; leino2020stolen; carlini2022membership; ye2022enhanced; zarifzadeh2024low; bertran2024scalable). This attack category directly maps to the differential privacy guarantee where two neighboring datasets differ in the presence of one training sample.
@@
A less studied category is attribute inference attacks (yeom2018privacy) where the adversary reconstructs a protected attribute given access to a partial record, of which label inference attacks are a special case (meta-labeldp; fekete2024auditing). A difficulty and common pitfall with such attacks is to properly account for the adversary’s baseline success, achieved by exploiting knowledge of the data distribution and correlations between the public and protected attributes (jayaraman2022are). The label inference attack of meta-labeldp uses canaries with random binary labels, which sets the adversary’s baseline accuracy to $0.5$ and allows Label DP to be audited via standard MIA analyses.
@@
Most auditing methods in the literature are “interventional,” as their privacy game involves manipulating the training dataset: MIAs require excluding a subset of the data from training (zanella2023bayesian; steinke2024privacy), whereas meta-labeldp modifies the training labels. Training-data interventions severely restrict the applications of these auditing methods and may degrade model performance if too many samples are withheld or out-of-distribution canaries are injected. Instead, our label inference attack can run entirely post-training. MIAs can be stated as observational privacy games if the challenger is able to sample fresh samples from the distribution (ye2022enhanced). However, obtaining new samples from the distribution (or from its close approximation), without affecting the training pipeline, remains an extremely challenging open problem (maini2024llm; duan2024do; meeus2025; blind_baselines25).
@@
kazmi2024panoramia propose an observational MIA that uses generative models such as generative adversarial networks (GANs) to obtain fresh samples. Differently from kazmi2024panoramia, we propose a more general observational auditing framework (with observational MIA and LIA as special cases), and provide theoretical lower bounds on the privacy guarantees for these general attacks. A challenge with using synthetic samples is to account for how the shift between the generative model and the true distribution boosts the success of the attacker. Counterfactual labels are much easier to generate than entire samples (models are trained to do precisely this), and we can assume that an adversary with limited computation resources cannot distinguish between a label from the distribution and a label predicted from the best model the adversary can train.
@@
The closest observational label inference attack to ours is the recent work of fekete2024auditing, which measures the label reconstruction advantage of the adversary with and without access to the model. This metric is not translated into a lower bound on $\epsilon$} in the differential privacy guarantee; in fact, such translation would be difficult because the adversary has a different baseline success (prior) for each sample. Additionally, this approach requires estimating probabilities of the mechanism’s output given a particular label, limiting its applicability to simple mechanisms like randomized response and random label aggregation. In contrast, our label inference attack is applicable to all mechanisms. It can audit label privacy in a statistically valid manner because it sets up a game where the baseline accuracy of the adversary is $0.5$ for all samples.
@@ 3 Preliminaries
In this work, we audit the simulation-based definition of DP, introduced by mahloujifar2025machine. It generalizes the traditional add/remove (or “leave-one-out”, or “zero-out”) notion of DP to support a privacy unit that is a subset of the sample’s attributes. The definition compares the distribution of a mechanism $M$ on dataset $D$ with that of a simulator that emulates the output of $M$ on $D$ without seeing the protected attributes of a record. This privacy definition fits nicely with our auditing framework, which is based on a proxy distribution (or simulator) that emulates training records. See Appendix 8 for further discussion of this definition.
Definition 1 (Simulation-based privacy for protected attributes (mahloujifar2025machine)).
@@
For the more advanced notion of $f$-DP (dong2019gaussian), a mechanism is f-SIM-DP with respect to $\mathrm{Sim}$ if $M(D)\approx_{f}\mathrm{Sim}(D^{\prime},x).$ We also say $M$ is $(\epsilon,\delta)$-SIM-DP (resp. f-SIM-DP) if there exists a simulator $\mathrm{Sim}$ for which (1) (resp. (1)) holds.
@@
An audit procedure takes the description of a mechanism $M$, a trade-off function $f$, a simulator $\mathrm{Sim}$ and decides whether the mechanism satisfies f-SIM-DP with respect to $\mathrm{Sim}$. We define it as a two-step process.
  • $\mathsf{Game}\colon\mathcal{M}\times\mathcal{S}\to\mathcal{O}$. The auditor runs a potentially randomized experiment/game using the description of mechanism $M\in\mathcal{M}$ and the simulator $\mathrm{Sim}$. The auditor receives the game output $o\in\mathcal{O}$.
  • $\mathsf{Evaluate}\colon\mathcal{O}\times\mathcal{F}\to\{0,1\}$. The output is 0 if the auditor rejects the hypothesis that $M$ satisfies f-SIM-DP based on evidence $o$, and $1$ otherwise.
@@ 4 Observational versus interventional privacy games
An observational privacy game considers the training dataset as a given. In contrast, an interventional privacy game interferes with the training data pipeline and the eventual dataset used for training. In this section, we formalize our observational privacy auditing framework. To that end, we introduce a generic attack game, generalizing swanberg2025, as Algorithm 1.
The privacy game occurs between two parties: a challenger and an adversary. A key difference from swanberg2025 is that we split the challenger algorithm into two stages: sampling the training data ($G_{\mathsf{data}}$) and sampling the additional outputs provided to the adversary ($G_{\mathsf{hint}}$). With this, we can separate observational games from interventional ones. Further, we distinguish between the training dataset $D$ and additional game artifacts $S$, which are used by the challenger to set up a stochastic game (e.g., sampling random bits).
Algorithm 1 Generic Attack Game (adapted from swanberg2025)
@@
The algorithm $G_{\mathsf{data}}$ determines the training dataset for $M$, obtained from a potential modification of the fixed training set $D$. For instance, in the one-run MIA (steinke2024privacy) the artifacts are $S=(S_{i}\sim\mathsf{Bernoulli}(0.5)\colon i\in[m])$, where $m$ is the number of canaries. The training dataset $\overline{D}$ is obtained from $D$ by including samples $x_{i}\in D$ where $S_{i}=1$ and excluding samples where $S_{i}=0$.
The role of $G_{\mathsf{hint}}$ is to collect additional information the challenger provides to the adversary, based on game artifacts, training data, and the output of the trained model $M(\overline{D})$. In the one-run MIA, $o_{2}$ is the vector of targets $x_{1},\dots,x_{m}$ from the training set $D$.
@@
Observational/interventional MIA. We have described how the (interventional) one-run MIA (steinke2024privacy) can be framed as Algorithm 1. We now present an observational one-run MIA (following ye2022enhanced).
@@
From the adversary’s perspective, the observational one-run MIA has the same distribution as the interventional MIA. For the challenger, the difference matters as the observational game does not alter the training pipeline. In this game, the source of counterfactual samples $x_{i}^{\prime}$ can be a distribution that approximates $\mathcal{D}$ sufficiently well, as is the case with the attack of kazmi2024panoramia who use generative models to sample $x_{i}^{\prime}$.